# Equitability and explosive synchronisation in multiplex and higher-order networks

### v1

@@ IX Conclusions
In this article, we give a full characterisation of all possible cluster synchronised solutions in terms of cluster independence, and equitability, for two very general systems of coupled dynamics (multiplex, and hypergraph). Namely, we show that cluster synchronisation, with respect to a partition $\mathcal{P}$, can only occur in either of these two cases:
@@
  • there is a non-trivial linear dependence relation between the coupling functions on trajectories (see Eqs.(11) and (22); or, if not,
@@
The first case depends on the particular algebraic properties of the coupling functions, while the second one is a structural property of the underlying pattern of interactions, and thus independent of the particular form of the coupling function.
Moreover, in the cluster independent, or equitable, case, we show that the correct notion of structural equitability is simultaneous equitability per layer (multiplex case, Eq. (10)) or per many-body interaction (hypergraph case, Eq. (21)), explaining why cluster synchronisation is much harder to achieve in general in multi-layer and higher-order coupled dynamical systems. Our theoretical findings are illustrated with numerical simulations for both multiplex networks and hypergraphs.
Although we cannot make any general claims about the stability of the cluster synchronised solutions, which will depend on the specific form of the coupling functions, and values of the coupling strength parameters, our results provide strong restrictions about the groupings of nodes that can independently synchronise for any choice of coupling function, based solely on the equitability of the underlying interaction network, multiplex, or hypergraph.
Finally, using the quotient dynamics, we also show that equitable solutions exists for any pre-determined clustering structure (even though we cannot say anything in general about their stability), paving the way for control applications in clustering engineering and control on arbitrary systems.
@@
Our results formalise and clarify the relationship between cluster synchronisation and equitability, including a new concept of dynamical stability, on networks, and higher-order networks, but several important open questions remain. These include fast and exhaustive algorithms to find equitable partitions in arbitrary multiplexes and hypergraphs; the realisation and ordering problem, that is, which equitable partitions and in which order they synchronise as we increase the coupling strength parameters (see Bayani et al. (2024) for the network case); the stability question, that is, finding general conditions that guarantee the stability and a synchronised solution, for instance from a quotient to a parent solution; and an extension to non-identical dynamical units such as general multi-layer networks and to other synchronisation types beyond identical synchronisation.

### v3

@@ III Conclusions
In this article, we are concerned with very general necessary and sufficient conditions for the existence of cluster synchronised solutions on two very general systems of coupled dynamics (multiplex, and hypergraph). We show that cluster synchronisation, with respect to a partition $\mathcal{P}$, can only occur in either of these two cases:
@@
  • there is a linear dependence relation between the coupling functions on trajectories (see Eqs.(11) and (22); or, if not,
@@
In other words, cluster synchronisation can occur without equitability (of the underlying cluster partition), but only if there are linear dependencies between the coupling functions. The first case (linear dependence) depends on the particular algebraic properties of the coupling functions, while the second one (equitability) is a structural property of the underlying pattern of interactions, and thus independent of the particular form of the coupling function.
Moreover, in the linear independent, or equitable, case, we show that the correct notion of external equitability is simultaneous equitability per layer (multiplex case, Eq. (10)) or per many-body interaction (hypergraph case, Eq. (21)), explaining why cluster synchronisation is much harder to achieve in general in multi-layer and higher-order coupled dynamical systems. Our theoretical findings are illustrated with numerical simulations for both multiplex networks and hypergraphs.
Although we cannot make any general claims about the stability of the cluster synchronised solutions, which will depend on the specific form of the coupling functions and values of the coupling strength parameters, our results provide strong restrictions about the groupings of nodes that can independently synchronise for any choice of coupling function, based solely on the equitability of the underlying interaction network, multiplex, or hypergraph.
Finally, using the quotient dynamics, we also show that equitable solutions exists for any pre-determined clustering structure, paving the way for control applications in clustering engineering and control on arbitrary systems.
There are limitations to our approach, for instance when there are linear dependencies between coupling functions either overall, or when restricted to some synchronised solutions: such generic solutions can exists without equitability of the underlying partition. (For a concrete example, inspired by the natural coupling in gambuzza2021stability , see the SI.) Again, such solutions are not ‘generic enough’ as they depend on linear dependencies of the coupling functions.
@@
Our results formalise and clarify the relationship between cluster synchronisation and equitability, but several important open questions remain. These include fast and exhaustive algorithms to find equitable partitions in arbitrary multiplexes and hypergraphs; the realisation and ordering problem, that is, which equitable partitions and in which order they synchronise as we increase the coupling strength parameters (see bayani2024transition for the network case); the stability question, that is, finding general conditions that guarantee the stability and a synchronised solution, for instance from a quotient to a parent solution; and an extension to non-identical dynamical units such as general multi-layer networks and to other synchronisation types beyond identical synchronisation. We hope our article will stimulate work on these, and other, directions.